Given that 43+x/810x = 210-2x/643x what is the value of x?


1
Expert's answer
2021-11-01T18:34:32-0400

"\frac{4^{3+x}}{8^{10x}} = \frac{2^{10-2x}} {64^{3x}}"


=> "4^{3+x} \u2022 64^{3x} = 8^{10x} \u2022 2^{10-2x}"

"2^{6+2x} \u2022 2^{18x} = 2^{30x} \u2022 2^{10-2x}"

"2^{6+2x+18x} = 2^{30x+10-2x}"


=> "6 + 2x +18x = 30x + 10 - 2x"

"20x - 28x = 10 - 6"

"-8x = 4"


=> "x = \frac{4}{-8} = \frac{-1}{2}"